# Multivariate generalized Pareto distributions along extreme directions

### v3

@@ 1 Introduction
The statistical description of extremes is a long-standing subject of research, with the ultimate goal of extrapolating beyond the range of the data and estimating probabilities of events more extreme than those observed in the past. The foundations of univariate extreme value theory o were established many decades ago through the derivation of the asymptotic distributions of suitably normalized block maxima and excesses over high thresholds. Modeling multivariate extreme events is a bigger challenge because, contrary to the univariate setting, the class of multivariate extreme-value distributions is infinite-dimensional. Over the last thirty years, numerous parametric extreme-value models have been proposed with the goal of capturing the tail dependence structure of a random $d$-dimensional vector $\boldsymbol{X}=(X_{1},\ldots,X_{d})$; see, e.g., Beirlant et al. (2004, Chapter 9) for an early overview or Cooley et al. (2010), Ballani and Schlather (2011) and Opitz (2013) for more recent examples. However, many such models are limited to the setting where all components of $\boldsymbol{X}$ can become large simultaneously. While this may be sufficient when the dimension $d$ of $\boldsymbol{X}$ is low, more often than not, one encounters the setting where some variables may be large while at the same time, the others are of smaller order. In finance, for example, characterizing groups of stocks within a portfolio that have the potential of suffering major losses simultaneously is of great importance. Managing flood risk requires identifying and modeling groups of gauging stations where extreme water discharges may occur simultaneously.
@@
Groups of variables that may become large without the others being so—in an asymptotic sense to be made precise below—will be called extreme directions. More specifically, if $J\subseteq\{1,\ldots,d\}$ is such that the components $(X_{j},j\in J)$ can be large simultaneously while the other components $(X_{j},j\in\{1,\ldots,d\}\setminus J)$ are small, then $J$ defines an extreme direction. The terminology is in line with a characterization of the support of the angular measure of the multivariate extreme-value distribution to which $\boldsymbol{X}$ is attracted. In the literature, such groups are also referred to as extreme clusters (Chiapino et al., 2020), extreme features (Jalalzai and Leluc, 2021), or concomitant extremes (Fomichov and Ivanovs, 2023). Many parametric extreme-value families are limited to the single extreme direction $J=\{1,\ldots,d\}$. The overarching goal of this paper is to explore the case when other extreme directions than $\{1,\ldots,d\}$ exist, and this in terms of several well-known objects in extreme-value theory, and multivariate generalized Pareto (mgp) distributions in particular.
@@
In the literature, tail dependence is described mathematically through various concepts, such as the exponent measure, the angular measure, and the stable tail dependence function, among others. When the focus is on modeling multivariate threshold exceedances, the mgp distribution is a natural candidate (Rootzén and Tajvidi, 2006). In the univariate case, peaks-over-threshold modeling with the generalized Pareto distribution is well-established (Davison and Smith, 1990), while in the multivariate set-up, statistical modeling is quite recent (Rootzén et al., 2018; Kiriliouk et al., 2018; Ho and Dombry, 2019; Engelke and Hitz, 2020; Hentschel et al., 2022). Asymptotically, the class of mgp distributions represents proper multivariate probability distributions on a region where at least one variable is large. Moreover, parametric mgp models can be constructed quite easily via the generators proposed in Kiriliouk et al. (2018).
@@
Learning extreme directions from data has become an active area of research. Goix et al. (2016, 2017) focus on the identification of extreme directions, providing a sparse representation of the support of the angular measure. Chiapino and Sabourin (2016) and Chiapino et al. (2019) study an alternative algorithm with better performance when the number of extreme directions is high. Jalalzai and Leluc (2021) introduce an approach based on empirical risk minimization. Chiapino et al. (2020) propose a Dirichlet mixture model for the angular measure of a heavy-tailed distribution whose support has been identified and apply it to anomaly detection. Simpson et al. (2020) exploit the concept of hidden regular variation to identify extreme directions. Finally, Meyer and Wintenberger (2021, 2023) frame extreme directions within what they call sparse regular variation.
@@
Our contribution differs from those above from several perspectives. First, our focus is on extreme directions in the context of multivariate threshold exceedances. Second, our aim is not to learn extreme directions, but rather to introduce parametric mgp models capable of generating extreme directions other than $\left\{{1,\ldots,d}\right\}$. For multivariate extreme-value distributions, the max-linear factor model (Einmahl et al., 2012) and the asymmetric logistic model (Tawn, 1990) already allow for this possibility. In the recent literature, only Simpson et al. (2020) and Chiapino et al. (2020) propose mixture models for multivariate extreme-value distributions. The power of our model construction resides in the fact that it is capable of representing mgp distributions with general extreme directions. Moreover, for specific parametric families, our model is straightforward to simulate from.
The paper is organized as follows. Section 2 recalls essential background on multivariate extreme-value theory. Section 3 defines and characterizes extreme directions in terms of the angular measure, the exponent measure, and the mgp distribution. We also compute the density of an mgp distribution with arbitrary extreme directions in terms of its generators, extending results in Kiriliouk et al. (2018). A model construction tool based on a mixture of mgp generators is presented in Section 4. Apart from enjoying a convenient interpretation as a smoothed max-linear model, the construction is able to reproduce any mgp distribution. For the sake of illustration, we work out two examples based on the logistic and the Hüsler–Reiss families. Moreover, the construction underlies a simulation algorithm for mgp distributions with non-standard extreme directions (Section 5). As a by-product, we revisit known simulation algorithms of standard mgp distributions in Dombry et al. (2016) and Engelke and Hitz (2020). Section 6 brings the paper to a close by outlining several avenues for future research. All proofs are deferred to the appendices.
@@ 2 Background
and let $\pi_{J}:\mathbb{A}_{J}\to\mathbb{R}^{\lvert J\rvert}$ be the projection $\boldsymbol{x}\mapsto\boldsymbol{x}_{J}$, where $\lvert J\rvert$ denotes the number of elements of the set $J$.
@@ 2.1 Multivariate extreme value distributions
For $d\geqslant 2$, the family of $d$-dimensional mev distributions no longer constitutes a finite-dimensional parametric family, in contrast to the univariate case. The reason is the dependence structure between the components of $G$, or equivalently, the extremal dependence between the components of $\boldsymbol{\xi}$. Many functions and measures that describe this dependence are discussed in the literature (see, e.g., Beirlant et al., 2004, Chapter 8). In this paper, three such objects will appear on the forefront: the exponent measure, the angular measure (de Haan and Resnick, 1977) and the stable tail dependence function (Drees and Huang, 1998).
To concentrate on the dependence structure of $G$, it is convenient to standardize its margins. The choice of common scale is a matter of convention and convenience. In this section, we will opt for the unit-Fréchet distribution, i.e., $\alpha_{j}=\gamma_{j}=\mu_{j}=1$ for all $j\in D$, in line with the theory as originally exposed in (de Haan and Resnick, 1977). In Section 2.2, however, we will choose the Gumbel distribution instead ($\alpha_{j}=1$ and $\gamma_{j}=\mu_{j}=0$ for all $j\in D$), in order to link up with the additive model formulation for the mgp in Rootzén et al. (2018). Switching from one standardization to another is an easy and sometimes illuminating exercise.
@@ 2.2 Multivariate generalized Pareto distributions
where $\boldsymbol{\zeta}$ is the vector of lower endpoints of $G_{1},\ldots,G_{d}$ and $\vee$ denotes the component-wise maximum. The conditioning event $\boldsymbol{\xi}\nleqslant\boldsymbol{b}_{n}$ is to be read as $\exists j\in D:\xi_{j}>b_{nj}$, that is, at least one variable within $\boldsymbol{\xi}$ exceeds a high threshold. We write $H=\operatorname{GP}(G)$. As in Rootzén et al. (2018), we consider the standardized case where $G$ has Gumbel margins, i.e., $\gamma_{j}=\mu_{j}=0$ and $\alpha_{j}=1$ for all $j\in D$. The distribution of an mgp random vector $\boldsymbol{Y}\sim H$ is then determined by the stdf $\ell$.
Recall the exponent measure $\mu$ of $G^{*}$. By Rootzén and Tajvidi (2006, Proposition 3.1), the law of $\mathrm{e}^{\boldsymbol{Y}}$ can be expressed in terms of $\mu$ on the collection of sets $\mathbb{A}=\left\{{[\boldsymbol{0},\boldsymbol{y}],\;\boldsymbol{y}\in\left[\boldsymbol{0},\boldsymbol{\infty}\right)}\right\}$, via
@@
In words, the law of $\mathrm{e}^{\boldsymbol{Y}}$ is equal to the restriction of $\mu$ to $\left\{\boldsymbol{x}\geqslant\boldsymbol{0}:\;\boldsymbol{x}\nleqslant\boldsymbol{1}\right\}$, normalized to a probability measure. Conversely, since $\mu\left({\left\{{\boldsymbol{x}\geqslant\boldsymbol{0}:\;x_{j}>1}\right\}}\right)=-\log G^{*}_{j}(1)=1$ by (2.6), we obtain the normalization constant via $\operatorname{\mathsf{P}}\left[{Y_{j}>0}\right]=1/\mu\left({\left\{{\boldsymbol{x}\geqslant\boldsymbol{0}:\;\boldsymbol{x}\nleqslant\boldsymbol{1}}\right\}}\right)$ for all $j\in D$, so that $\mu$ can be reconstructed from $\boldsymbol{Y}$ by
@@
Rootzén et al. (2018, Section 4) propose three stochastic representations for $H$. In view of their importance to the construction device and the simulation algorithm to follow, we review these representations in some detail.
@@
Let $\boldsymbol{S}$ be a random vector in $[-\infty,0]^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{S})=0]=1$, $\operatorname{\mathsf{P}}[S_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{S_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{S_{d}}]$. Then $\boldsymbol{S}$ is said to be the spectral random vector (Rootzén et al., 2018, Theorem 6) associated to the mgp $H$ if
$\ell(\boldsymbol{y})=\operatorname{\mathsf{E}}[\max\{\boldsymbol{y}\mathrm{e}^{\boldsymbol{S}}\}]/\operatorname{\mathsf{E}}[\mathrm{e}^{S_{1}}],\qquad\boldsymbol{y}\in\left[\boldsymbol{0},\boldsymbol{\infty}\right).$ (2.10)
@@
with $E=\max(\boldsymbol{Y})$ a unit exponential random variable independent of the spectral random vector $\boldsymbol{S}=\boldsymbol{Y}-\max(\boldsymbol{Y})$; see Rootzén et al. (2018, Theorem 7).
@@
Let $\boldsymbol{T}$ be a random vector in $\left[-\infty,\infty\right)^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{T})>-\infty]=1$, $\operatorname{\mathsf{P}}[T_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{T_{1}-\max(\boldsymbol{T})}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{T_{d}-\max(\boldsymbol{T})}]$. The random vector $\boldsymbol{T}$ is called a $\boldsymbol{T}$-generator of $H$ (Rootzén et al., 2018, Proposition 8), if we have the equality in distribution
@@
Let $\boldsymbol{U}$ be a random vector on $[-\infty,\infty)^{d}$ such that $\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{U_{d}}]$ and $0<\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]<\infty$. The random vector $\boldsymbol{U}$ is said to be an $\boldsymbol{U}$-generator of $H$ (Rootzén et al., 2018, Proposition 9) if
@@
If $\boldsymbol{U}$ is a generator of $\boldsymbol{Y}\sim H$, then, for non-empty $J\subset D$, the random vector $\boldsymbol{U}_{J}$ is a generator of $\boldsymbol{Y}_{J}\mid\max(\boldsymbol{Y}_{J})>0$, which thus follows a $|J|$-variate mgp (Rootzén et al., 2018, Proposition 18). In Proposition 5.2 below, we establish a link between $\boldsymbol{U}$ and the extremal functions in Dombry et al. (2016). Equation (2.14) represents the stdf $\ell$ as a $D$-norm with generator $\boldsymbol{U}$ in the sense of Falk (2019). In Kiriliouk et al. (2018, Section 7), common parametric families of multivariate extreme value models are described in terms of their $\boldsymbol{U}$-generator.
@@
Note that a spectral random vector $\boldsymbol{S}$ is also a $\boldsymbol{T}$-generator, since $\max(\boldsymbol{S})=0$ and thus $\operatorname{GP}_{\boldsymbol{T}}(\mathcal{L}(\boldsymbol{S}))=\operatorname{GP}_{\boldsymbol{S}}(\mathcal{L}(\boldsymbol{S}))$. On the other hand, the representation via $\boldsymbol{U}$ is of another nature: if $\boldsymbol{V}$ is a random vector whose distribution satisfies both the constraints to be a $\boldsymbol{T}$-generator and a $\boldsymbol{U}$-generator, then the mgp distributions $\operatorname{GP}_{\boldsymbol{T}}(\mathcal{L}(\boldsymbol{V}))$ and $\operatorname{GP}_{\boldsymbol{U}}(\mathcal{L}(\boldsymbol{V}))$ are different in general; see Kiriliouk et al. (2018, Section 7) for examples.

### v4

@@ 1 Introduction
The statistical description of extremes is a long-standing subject of research, with the ultimate goal of extrapolating beyond the range of the data and estimating probabilities of events more extreme than those observed in the past. The foundations of univariate extreme value theory were established many decades ago through the derivation of the asymptotic distributions of suitably normalized block maxima and excesses over high thresholds. Modeling multivariate extreme events is a bigger challenge because, contrary to the univariate setting, the class of multivariate extreme-value distributions is infinite-dimensional. Over the last thirty years, numerous parametric extreme-value models have been proposed with the goal of capturing the tail dependence structure of a random $d$-dimensional vector $\boldsymbol{X}=(X_{1},\ldots,X_{d})$; see, e.g., Beirlant et al. [2, Chapter 9] for an early overview or Cooley et al. [9], Ballani and Schlather [1] and Opitz [33] for more recent examples. However, many such models are limited to the setting where all components of $\boldsymbol{X}$ can become large simultaneously. While this may be sufficient when the dimension $d$ of $\boldsymbol{X}$ is low, more often than not, one encounters the setting where some variables may be large while at the same time, the others are of smaller order. In finance, for example, characterizing groups of stocks within a portfolio that have the potential of suffering major losses simultaneously is of great importance. Managing flood risk requires identifying and modeling groups of gauging stations where extreme water discharges may occur simultaneously.
@@
Groups of variables that may become large without the others being so—in an asymptotic sense to be made precise below—will be called extreme directions. More specifically, if $J\subseteq\{1,\ldots,d\}$ is such that the components $(X_{j},j\in J)$ can be large simultaneously while the other components $(X_{j},j\in\{1,\ldots,d\}\setminus J)$ are small, then $J$ defines an extreme direction. The terminology is in line with a characterization of the support of the angular measure of the multivariate extreme-value distribution to which $\boldsymbol{X}$ is attracted. In the literature, such groups are also referred to as extreme clusters [6], extreme features [27], or concomitant extremes [20]. Many parametric extreme-value families are limited to the single extreme direction $J=\{1,\ldots,d\}$. The overarching goal of this paper is to explore the case when other extreme directions than $\{1,\ldots,d\}$ exist, and this in terms of several well-known objects in extreme-value theory, and multivariate generalized Pareto (mgp) distributions in particular.
@@
In the literature, tail dependence is described mathematically through various concepts, such as the exponent measure, the angular measure, and the stable tail dependence function, among others. When the focus is on modeling multivariate threshold exceedances, the mgp distribution is a natural candidate [34]. In the univariate case, peaks-over-threshold modeling with the generalized Pareto distribution is well-established [10], while in the multivariate set-up, statistical modeling is quite recent [35, 29, 24, 16, 23]. Asymptotically, the class of mgp distributions represents proper multivariate probability distributions on a region where at least one variable is large. Moreover, parametric mgp models can be constructed quite easily via the generators proposed in Kiriliouk et al. [29].
@@
Learning extreme directions from data has become an active area of research. Goix et al. [21, 22] focus on the identification of extreme directions, providing a sparse representation of the support of the angular measure. Chiapino and Sabourin [4] and Chiapino et al. [5] study an alternative algorithm with better performance when the number of extreme directions is high. Jalalzai and Leluc [27] introduce an approach based on empirical risk minimization. Chiapino et al. [6] propose a Dirichlet mixture model for the angular measure of a heavy-tailed distribution whose support has been identified and apply it to anomaly detection. Simpson et al. [39] exploit the concept of hidden regular variation to identify extreme directions. Finally, Meyer and Wintenberger [31, 32] frame extreme directions within what they call sparse regular variation.
@@
Our contribution differs from those above from several perspectives. First, our focus is on extreme directions in the context of multivariate threshold exceedances. Second, our aim is not to learn extreme directions, but rather to introduce parametric mgp models capable of generating extreme directions other than $\left\{{1,\ldots,d}\right\}$. For multivariate extreme-value distributions, the max-linear factor model [15] and the asymmetric logistic model [41] already allow for this possibility. In the recent literature, only Simpson et al. [39] and Chiapino et al. [6] propose mixture models for multivariate extreme-value distributions. The power of our model construction resides in the fact that it is capable of representing mgp distributions with general extreme directions. Moreover, for specific parametric families, our model is straightforward to simulate from.
The paper is organized as follows. Section 2 recalls essential background on multivariate extreme-value theory. Section 3 defines and characterizes extreme directions in terms of the angular measure, the exponent measure, and the mgp distribution. We also compute the density of an mgp distribution with arbitrary extreme directions in terms of its generators, extending results in Kiriliouk et al. [29]. A model construction tool based on a mixture of mgp generators is presented in Section 4. Apart from enjoying a convenient interpretation as a smoothed max-linear model, the construction is able to reproduce any mgp distribution. For the sake of illustration, we work out two examples based on the logistic and the Hüsler–Reiss families. Moreover, the construction underlies a simulation algorithm for mgp distributions with non-standard extreme directions (Section 5). As a by-product, we revisit known simulation algorithms in Dombry et al. [12] and Engelke and Hitz [16] that can be used to simulate mgp random vectors. Section 6 brings the paper to a close by outlining several avenues for future research. All proofs are deferred to the appendices.
@@ 2 Background
and let $\pi_{J}:\mathbb{A}_{J}\to\mathbb{R}^{J}$ be the projection $\boldsymbol{x}\mapsto\boldsymbol{x}_{J}$ and with $\mathbb{R}^{J}$ the set of real vectors $(x_{j})_{j\in J}$ indexed by $J$. Clearly, $\mathbb{R}^{J}$ is isomorphic to $\mathbb{R}^{|J|}$, with $\lvert J\rvert$ the number of elements of $J$, but it will be convenient to keep track of the index set explicitly, not only its cardinality.
@@ 2.1 Multivariate extreme value distributions
For $d\geqslant 2$, the family of $d$-dimensional mev distributions no longer constitutes a finite-dimensional parametric family, in contrast to the univariate case. This is due to the nature of the dependence structure between the components of $G$, or equivalently, the extremal dependence between the components of $\boldsymbol{\xi}$. Many functions and measures that describe this dependence are discussed in the literature [see, e.g., 2, Chapter 8]. In this paper, three such objects will appear on the forefront: the exponent measure, the angular measure [11] and the stable tail dependence function [13].
To concentrate on the dependence structure of $G$, it is convenient to standardize its margins. The choice of common scale is a matter of convention and convenience. In this section, we will opt for the unit-Fréchet distribution, i.e., $\alpha_{j}=\gamma_{j}=\mu_{j}=1$ for all $j\in D$, in line with the theory as originally exposed in [11]. In Section 2.2, however, we will choose the Gumbel distribution instead ($\alpha_{j}=1$ and $\gamma_{j}=\mu_{j}=0$ for all $j\in D$), in order to link up with the additive model formulation for the mgp in Rootzén et al. [36]. Switching between different marginal scales is an easy and sometimes illuminating exercise.
@@ 2.2 Multivariate generalized Pareto distributions
where $\boldsymbol{\zeta}$ is the vector of lower endpoints of $G_{1},\ldots,G_{d}$ and $\vee$ denotes the component-wise maximum. The conditioning event $\boldsymbol{\xi}\nleqslant\boldsymbol{b}_{n}$ is to be read as $\exists j\in D:\xi_{j}>b_{nj}$, that is, at least one variable within $\boldsymbol{\xi}$ exceeds a high threshold. We write $H=\operatorname{GP}(G)$. As in Rootzén et al. [36], we consider the standardized case where $G$ has Gumbel margins, i.e., $\gamma_{j}=\mu_{j}=0$ and $\alpha_{j}=1$ for all $j\in D$. The distribution of an mgp random vector $\boldsymbol{Y}\sim H$ is then determined by the stdf $\ell$.
Recall the exponent measure $\mu$ of $G^{*}$. By Rootzén and Tajvidi [34, Proposition 3.1], the law of $\mathrm{e}^{\boldsymbol{Y}}$ can be expressed in terms of $\mu$ on the collection of sets $\mathbb{A}=\left\{{[\boldsymbol{0},\boldsymbol{y}],\;\boldsymbol{y}\in\left[\boldsymbol{0},\boldsymbol{\infty}\right)}\right\}$, via
@@
In words, the law of $\mathrm{e}^{\boldsymbol{Y}}$ is equal to the restriction of $\mu$ to $\left\{\boldsymbol{x}\geqslant\boldsymbol{0}:\;\boldsymbol{x}\nleqslant\boldsymbol{1}\right\}$, normalized to a probability measure. Conversely, since $\mu\left({\left\{{\boldsymbol{x}\geqslant\boldsymbol{0}:\;x_{j}>1}\right\}}\right)=-\ln G^{*}_{j}(1)=1$ by (2.6), we obtain the normalization constant via $\operatorname{\mathsf{P}}\left[{Y_{j}>0}\right]=1/\mu\left({\left\{{\boldsymbol{x}\geqslant\boldsymbol{0}:\;\boldsymbol{x}\nleqslant\boldsymbol{1}}\right\}}\right)$ for all $j\in D$, so that $\mu$ can be reconstructed from $\boldsymbol{Y}$ by
@@
Rootzén et al. [36, Section 4] propose three stochastic representations for $H$. In view of their importance to the construction device and the simulation algorithm to follow, we review these representations in some detail.
@@
Let $\boldsymbol{S}$ be a random vector in $[-\infty,0]^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{S})=0]=1$, $\operatorname{\mathsf{P}}[S_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{S_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{S_{d}}]$. Then $\boldsymbol{S}$ is said to be the spectral random vector [36, Theorem 6] associated to the mgp $H$ if
$\ell(\boldsymbol{y})=\operatorname{\mathsf{E}}[\max\left({\boldsymbol{y}\mathrm{e}^{\boldsymbol{S}}}\right)]/\operatorname{\mathsf{E}}[\mathrm{e}^{S_{1}}],\qquad\boldsymbol{y}\in\left[\boldsymbol{0},\boldsymbol{\infty}\right).$ (2.10)
@@
with $E=\max(\boldsymbol{Y})$ a unit exponential random variable independent of the spectral random vector $\boldsymbol{S}=\boldsymbol{Y}-\max(\boldsymbol{Y})$; see Rootzén et al. [36, Theorem 7].
@@
Let $\boldsymbol{T}$ be a random vector in $\left[-\infty,\infty\right)^{d}$ such that $\operatorname{\mathsf{P}}[\max(\boldsymbol{T})>-\infty]=1$, $\operatorname{\mathsf{P}}[T_{j}>-\infty]>0$ for all $j\in D$ and $\operatorname{\mathsf{E}}[\mathrm{e}^{T_{1}-\max(\boldsymbol{T})}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{T_{d}-\max(\boldsymbol{T})}]$. The random vector $\boldsymbol{T}$ is called a $\boldsymbol{T}$-generator of $H$ [36, Proposition 8], if we have the equality in distribution
@@
Let $\boldsymbol{U}$ be a random vector on $[-\infty,\infty)^{d}$ such that $\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]=\ldots=\operatorname{\mathsf{E}}[\mathrm{e}^{U_{d}}]$ and $0<\operatorname{\mathsf{E}}[\mathrm{e}^{U_{1}}]<\infty$. The random vector $\boldsymbol{U}$ is said to be a $\boldsymbol{U}$-generator of $H$ [36, Proposition 9] if
@@
If $\boldsymbol{U}$ is a generator of $\boldsymbol{Y}\sim H$, then, for non-empty $J\subset D$, the random vector $\boldsymbol{U}_{J}$ is a generator of $\boldsymbol{Y}_{J}\mid\max(\boldsymbol{Y}_{J})>0$, which thus follows a $|J|$-variate mgp [36, Proposition 18]. In Proposition 5.2 below, we establish a link between $\boldsymbol{U}$ and the extremal functions in Dombry et al. [12]. Equation (2.14) represents the stdf $\ell$ as a $D$-norm with generator $\boldsymbol{U}$ in the sense of Falk [19]. In Kiriliouk et al. [29, Section 7], common parametric families of multivariate extreme value models are described in terms of their $\boldsymbol{U}$-generator.
@@
Note that a spectral random vector $\boldsymbol{S}$ is also a $\boldsymbol{T}$-generator, since $\max(\boldsymbol{S})=0$ and thus $\operatorname{GP}_{\boldsymbol{T}}(\mathcal{L}(\boldsymbol{S}))=\operatorname{GP}_{\boldsymbol{S}}(\mathcal{L}(\boldsymbol{S}))$. On the other hand, the representation via $\boldsymbol{U}$ is of another nature: if $\boldsymbol{V}$ is a random vector whose distribution satisfies both the constraints to be a $\boldsymbol{T}$-generator and a $\boldsymbol{U}$-generator, then the mgp distributions $\operatorname{GP}_{\boldsymbol{T}}(\mathcal{L}(\boldsymbol{V}))$ and $\operatorname{GP}_{\boldsymbol{U}}(\mathcal{L}(\boldsymbol{V}))$ are different in general; see Kiriliouk et al. [29, Section 7] for examples.